OpenStudy (anonymous):

log(1/100).. and how?

OpenStudy (anonymous):

log(10^(-2))= -2log(10)

OpenStudy (anonymous):

log(1/100) = log(100)^-1 = log(10)^-2 = -2log(10) = -2

OpenStudy (anonymous):

log is usually base 10 so they should cancel out with the 10 inside so final answer should b -2 as nayeaddo put

OpenStudy (anonymous):

great thanks guys!

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

You can also remember the important rule that log a - log b = log(a/b)

OpenStudy (anonymous):

So this is log(1) - log(100) = 0 - 2 = -2

OpenStudy (radar):

Good point mtbender

OpenStudy (anonymous):

tru u can do that. u should have all the rules written down sumwhere as a reference. my teacher made us do that n it helped us to remember them